# Improving Medical Multi-modal Contrastive Learning with Expert Annotations

### v1

@@ 4.1.3 Training
[…]
@@ 4.2 Zero-shot Image Classification
Following CLIP(Radford et al., 2021), our zero-shot classification method categorizes images into predefined classes without direct finetuning, thus relying on the quality of embeddings generated during pretraining. To formulate the embedding of each class label, we first generate descriptive prompts to obtain a list of text embeddings corresponding to the label using the text encoder (Huang et al., 2021; Wang et al., 2022; You et al., 2023). The mean of these embeddings is taken as the representation of the label. For each image, classification is then performed by matching the image embedding with its closest label embedding through cosine similarity. More details of the prompts used are provided in the Supplement.
Table 1 illustrates the zero-shot classification performance on CheXpert 5x200, MIMIC 5x200, RSNA, and CXR 14x100 datasets. The results, based on macro-averaged F1 scores from three random initializations, highlight eCLIP variants’ superior performance over the base models across all datasets. While the $m^{3}$-mixup excels in MIMIC for certain architectures, eCLIP variants show broader generalization. Notably, eCLIP’s advantages are more pronounced in multi-class scenarios (CheXpert, MIMIC and CXR14) compared to binary classification on RSNA.
@@ 4.3 Sample Efficiency
Sample efficiency measures how well a model learns from limited amount of training data. eCLIP improves this efficiency by using expert annotated images to form new positive and negative pairs, aiming to improve the quality of the learned embeddings. To test this, we first looked at zero-shot classification performance, adjusting the number of training batches available for pretraining. Results, shown in Fig. 4 (top row), reveal that eCLIP is more sample efficient across MIMIC 5x200, CheXpert 5x200 and CXR 14x100 datasets compared to the base model. Additionally, by applying supervised fine-tuning (SFT) with a linear probe on class-imbalanced datasets – CXR-8, RSNA and OpenI-5 (Section 4.1.2) – eCLIP demonstrates stronger performance in multi-label classification tasks, CXR-8 and OpenI-5 and remains competitive in binary classification for RSNA. This is shown Fig. 4 (bottom row) where we plot the ROC AUC scores against different training sample sizes for linear probing. These findings highlight eCLIP’s ability to effectively learn from fewer samples.
@@ 4.4 Text retrieval and retrieval augmented generation (RAG)
${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ 2.9 7.5 10.3
[…]
@@
To compare eCLIP’s cross-modal functionality with that of CLIP we focused on text retrieval task using the Open-I dataset, which consists of pairs of X-rays and radiology reports. We used the FAISS vector database (Douze et al., 2024) to index the text embeddings generated by the text encoder. For a given X-ray image $I_{i}$, we then retrieve the closest text reports from the database based on the cosine similarity in the embedding space, $min_{j}(v_{i}\cdot t_{j})$. Results in Table 4 compare the performance of eCLIP against CLIP in text retrieval measured in Recall@1, 5, and 10. The performance of eCLIP indicates a notable improvement in its embedding quality. Note that our evaluation followed a strict criterion for recall computation, where a retrieval was counted as successful only if the exact correct report was identified. While more nuanced measures based on semantic similarity could be employed (You et al., 2023), we opted this approach to maintain a clear and simple evaluation framework.
Next we extend our analysis from retrieval to report generation using a frozen Large Language Model (LLM), Mistral 7B Instruct v2 (Jiang et al., 2023), aiming to generate radiology reports through Retrieval Augmented Generation (RAG). This setup tests the CLIP model’s capacity to retrieve texts which can be used to prompt an LLM to generate a report without finetuning on medical data. First we randomly selected 389 samples from the Open-I dataset for testing and utilized the FAISS database to index the reports from the remaining samples (i.e., training set). Given a test image we retrieve five closest reports from the training set and use them in the prompt for the LLM to generate a report for the test image. The eCLIP variant showed a small but consistent improvement over the base model in generating reports, as indicated in Table 4. A comparative analysis of generated report versus ground truth shown in Table 5, with discrepancies marked, further validates the effectiveness of eCLIP’s embeddings in supporting complex cross-modal tasks. Additional details, including LLM prompts and generated report samples are available in the Supplement.
@@ 4.6 Ablation Study
Our ablation study with the Swin Tiny encoder shows the impact of key components in our eCLIP model: multi-headed attention (MHA) layer for heatmap procesing, curriculum learning for phased introduction of expert annotations, mixup augmentation to compensate for limited number of export annotated data and priming of heatmap processor during initial training phase. Results shown in Table 4 reveal that the MHA-based heatmap processor improves zero-shot classification performance on CheXpert 5x200 and CXR 14x100 datasets compared to basic methods like direct application of heatmaps as mask ($\odot Mask)$ or using a CNN encoder. We note a significant performance drop with randomly generated heatmaps versus expert eye-gaze heatmaps. This highlights that while our methodological improvements contribute to the performance gains, the integration of meaningful, expert-derived signals is essential for achieving optimal results.
@@ 5 Conclusion
We introduce eCLIP, an adaptation of CLIP, demonstrating the integration of radiologist eye-tracking heatmap to overcome challenges faced in multi-modal contrastive learning. This study highlights the impact of integrating these high-quality expert annotations on improving the quality of learned embeddings and assess its influence on sample efficiency and cross-modal retrieval tasks. An important future research direction would be extending this approach to include expert annotations from the text modality (e.g., by adapting SimCSE (Gao et al., 2021)) and to leverage the temporal dynamics of eye-tracking data by aligning the sequential frames with the corresponding report snippets.
@@ 5.0.1 Limitations
Our study is limited by the small size of expert annotated data and thus does not comprehensively analyze the impact of size or distribution of expert annotations across different abnormalities. eCLIP also incurs extra computational costs during training due to the additional forward pass required for processing expert images in the warmup and cool-down phases. Additionally, the clinical relevance of generated radiology reports has not been validated by medical experts, relying instead on standard metrics known for potential biases and inaccuracies in reflecting clinical accuracy(Yu et al., 2023).

### v3

@@ 4.1.3 Training
We employed CLIP pretraining on the MIMIC-CXR dataset, utilizing a subset with roughly 1000 images with expert eye-gaze heatmap annotations, while validation and all other downstream evaluations proceed without these annotations. Our model architecture includes Swin Tiny, following recent studies[54, 59], alongside Vision Transformer (ViT) Small and Base, with image encoders pretrained on ImageNet and Clinical BERT [1] with max length of sequences set to 256 as the text encoder. We cropped images to $(224,224)$ using random resized crop augmentation and turned off all other image augmentations. Pretraining utilized 8 AMD MI250X GPUs, maintaining an effective batch size of 512 for 10,000 steps. The learning rate was $1e^{-4}$ for standard CLIP, increased to $2e^{-4}$ for eCLIP variants, with cosine annealing plus a linear warmup for the first 10% of iterations, weight decay of $1e^{-3}$, and learnable temperature parameter in the contrastive loss initialized at 0.07. Models for $m^{3}$-mixup and eCLIPP are initialized with weights from CLIP pretraining and further finetuned for 1,000 iterations with a learning rate of $1e^{-5}$. Detailed setup information is available in the Supplement.
@@ 4.2 Zero-shot Image Classification
Following CLIP[36], our zero-shot classification method categorizes images into predefined classes without direct finetuning, thus relying on the quality of embeddings generated during pretraining. To formulate the embedding of each class label, we first generate descriptive prompts to obtain a list of text embeddings corresponding to the label using the text encoder [15, 54, 59]. The mean of these embeddings is taken as the representation of the label. For each image, classification is then performed by matching the image embedding with its closest label embedding through cosine similarity. More details of the prompts used are provided in the Supplement.
Tab. 1 illustrates the zero-shot classification performance on CheXpert 5x200, MIMIC 5x200, RSNA, and CXR 14x100 datasets. The results, based on macro-averaged F1 scores from three random initializations, highlight eCLIP variants’ superior performance over the base models across all datasets. While the $m^{3}$-mixup excels in MIMIC for certain architectures, eCLIP variants show broader generalization. Notably, eCLIP’s advantages are more pronounced in multi-class scenarios (CheXpert, MIMIC and CXR14) compared to binary classification on RSNA.
@@ 4.3 Sample Efficiency
Sample efficiency measures how well a model learns from limited amount of training data. eCLIP improves this efficiency by using expert annotated images to form new positive and negative pairs, aiming to improve the quality of the learned embeddings. To test this, we first looked at zero-shot classification performance, adjusting the number of training batches available for pretraining. Results, shown in Fig. 4 (top row), reveal that eCLIP is more sample efficient across MIMIC 5x200, CheXpert 5x200 and CXR 14x100 datasets compared to the base model. Additionally, by applying supervised fine-tuning (SFT) with a linear probe on class-imbalanced datasets – CXR-8, RSNA and OpenI-5 (Sec. 4.1.2) – eCLIP demonstrates stronger performance in multi-label classification tasks, CXR-8 and OpenI-5 and remains competitive in binary classification for RSNA. This is shown Fig. 4 (bottom row) where we plot the ROC AUC scores against different training sample sizes for linear probing. These findings highlight eCLIP’s ability to effectively learn from fewer samples.
@@ 4.4 Text retrieval and retrieval augmented generation (RAG)
   ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ 3.9 9.4 13.2
Table 3: Performance on report generation with Open-I dataset. We report the BLEU-2 score (BL-2), BERT recall score[64] (B-R), Cosine similarity between the sentence embeddings of the generated and ground-truth report for MPNet [42] Sentence Transformer model [38] (Semb), and for the CheXBERT model[41] (CBemb)
Model BL-2 B-R Semb CBemb
@@
To compare eCLIP’s cross-modal functionality with that of CLIP we focused on text retrieval task using the Open-I dataset, which consists of pairs of X-rays and radiology reports. We used the FAISS vector database [8] to index the text embeddings generated by the text encoder. For a given X-ray image $I_{i}$, we then retrieve the closest text reports from the database based on the cosine similarity in the embedding space, $min_{j}(v_{i}\cdot t_{j})$. Results in Tab. 4 compare the performance of eCLIP against CLIP in text retrieval measured in Recall@1, 5, and 10. The performance of eCLIP indicates a notable improvement in its embedding quality. Note that our evaluation followed a strict criterion for recall computation, where a retrieval was counted as successful only if the exact correct report was identified. While more nuanced measures based on semantic similarity could be employed [59], we opted this approach to maintain a clear and simple evaluation framework.
Next we extend our analysis from retrieval to report generation using a frozen Large Language Model (LLM), Mistral 7B Instruct v2 [19], aiming to generate radiology reports through Retrieval Augmented Generation (RAG). This setup tests the CLIP model’s capacity to retrieve texts which can be used to prompt an LLM to generate a report without finetuning on medical data. First we randomly selected 389 samples from the Open-I dataset for testing and utilized the FAISS database to index the reports from the remaining samples (i.e., training set). Given a test image we retrieve five closest reports from the training set and use them in the prompt for the LLM to generate a report for the test image. The eCLIP variant showed a small but consistent improvement over the base model in generating reports, as indicated in Tab. 4. A comparative analysis of generated report versus ground truth shown in Tab. 5, with discrepancies marked, further validates the effectiveness of eCLIP’s embeddings in supporting complex cross-modal tasks. Additional details, including LLM prompts and generated report samples are available in the Supplement.
@@ 4.6 Ablation Study
Our ablation study with the Swin Tiny encoder shows the impact of key components in our eCLIP model: multi-headed attention (MHA) layer for heatmap procesing, curriculum learning for phased introduction of expert annotations, mixup augmentation to compensate for limited number of export annotated data and priming of heatmap processor during initial training phase. Results shown in Tab. 4 reveal that the MHA-based heatmap processor improves zero-shot classification performance on CheXpert 5x200 and CXR 14x100 datasets compared to basic methods like direct application of heatmaps as mask ($\odot Mask)$ or using a CNN encoder. We note a significant performance drop with randomly generated heatmaps versus expert eye-gaze heatmaps. This highlights that while our methodological improvements contribute to the performance gains, the integration of meaningful, expert-derived signals is essential for achieving optimal results.
@@ 5 Conclusion
We introduce eCLIP, an adaptation of CLIP, demonstrating the integration of radiologist eye-tracking heatmap to overcome challenges faced in multi-modal contrastive learning. This study highlights the impact of integrating these high-quality expert annotations on improving the quality of learned embeddings and assess its influence on sample efficiency and cross-modal retrieval tasks. An important future research direction would be extending this approach to include expert annotations from the text modality (e.g., by adapting SimCSE [9]) and to leverage the temporal dynamics of eye-tracking data by aligning the sequential frames with the corresponding report snippets.
@@ 5.0.1 Limitations
Our study is limited by the small size of expert annotated data and thus does not comprehensively analyze the impact of size or distribution of expert annotations across different abnormalities. eCLIP also incurs extra computational costs during training due to the additional forward pass required for processing expert images in the warmup and cool-down phases. Additionally, the clinical relevance of generated radiology reports has not been validated by medical experts, relying instead on standard metrics known for potential biases and inaccuracies in reflecting clinical accuracy[60].
Acknowledgements
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI, and grants 352986, 358246) and EU (H2020 grant 101016775 and NextGenerationEU). We acknowledge the computational resources provided by Aalto Science-IT project. We acknowledge CSC for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through Finland.
Supplement
[…]
Appendix 0.A Detailed Experiment Setup
0.A.1 Pretraining with Expert Annotations
For pretraining both CLIP and eCLIP models, we utilize the MIMIC-CXR dataset. Expert annotations, in the form of heatmaps, are derived from a subset of MIMIC-CXR, the EGD-CXR dataset [22], which comprises of 1080 samples. We employ the author’s official preprocessing code to convert the eye-tracking fixation data into heatmaps. The data tuple (image, text, heatmap) is then used for training with contrastive learning. We employ two dataloader: one for the main dataset without heatmap (“main loader”) and another for the subset with expert heatmaps (“expert loader”). In each training iteration, one batch is fetched from each loader; the CLIP model processed only the main batch, while the eCLIP model has the flexibility to use either just the main batch or both batches. LABEL:lst:pseudo-code shows the Pytorch-like pseudocode for the eCLIP model.
The utilization of the expert batch in eCLIP is determined by a curriculum probability, initially set to zero during the cold start phase. This probability linearly increases to $p_{max}$ during the warmup phase, then linearly decreases to $p_{min}$ during the cool-down phase, where it remains for the remainder of the training process. $p_{max}$ was set to 0.5 for all experiments, while $p_{min}$ was set to 0.05 for the ViT Base model and to 0.1 for all other models.
1def heatmap_processor(image, heatmap):
2 # reshapes and transposes have been omitted for brevity
3 pathches = patchify(image)
4 heatmap_patches = patchify(heatmap * image)
5 processed_patches = multi_head_attention(
6 q=heatmap_patches, k=patches, v=patches)
7 reconstructed_image = unpatchify(processed_patches)
8 return reconstructed_image
9
10class ExpertClipImageEncoder:
11 self.base = ViT()
12 self.projector = ProjectionBlock()
13
14 def forward(image, heatmap=None):
15 mse_loss = None
16 B, C, H, W = image.size()
17 # priming
18 reconstructed_image = heatmap_processor(
19 image, torch.ones((B, 1, H, W)))
20 mse_loss = mse_loss_fn(image, reconstructed_image)
21
22 image_features = self.base(image)
23 # project to shared embedding dimension
24 image_embed = self.projector(image_features)
25
26 if heatmap is not None:
27 expert_image = heatmap_processor(image, heatmap)
28 # mixup augmentation
29 lambda_ = beta(alpha=0.3).sample()
30 expert_image_m = mixup(image, expert_image, lambda_)
31
32 expert_image_features = self.base(expert_image_m)
33 expert_image_embed = self.projector(expert_image_features)
34
35 return image_embed, expert_image_embed, mse_loss
36
37# include expert batch based on curriculum learning probability
38# curriculum_prob is varied between cpmin to cpmax (typically, 0.1 & 0.5)
39use_expert = np.random.rand() < curriculum_prob
40
41# compute clip loss
42clip_loss = clip_loss_fn(
43 concat([image_embed, expert_image_embed]) if use_expert else image_embed,
44 concat([text_embed, text_embed]) if use_expert else text_embed
45)
Listing 1: PyTorch-like pseudocode for eCLIP implementation
0.A.2 $m^{2}$-mixup vs $m^{3}$-mixup
We illustrate $m^{2}$-mixup in Fig. 3, where embeddings from image and text domains are mixed to mitigate the modality gap, as proposed by Oh et al. [33]. Oh et al. [33] further introduce $m^{3}$-mixup, which combines $m^{2}$-mixup with corresponding unimodal mixups. Specifically,
$\mathcal{L}_{m^{3}}=\mathcal{L}_{\text{CLIP}}+\mathcal{L}_{m^{2}}+\mathcal{L}_{uni}$
For further details, please refer to Oh et al. [33].
0.A.3 Linear Probe Experiments
In our linear probe experiments, we utilize the CLIP and eCLIP with Swin Tiny as the image encoder following other recent similar works [54, 59]. The pretrained model’s image encoder is entirely frozen and we append a linear layer for classification. This layer’s output dimension is set to 1 for the Pneumonia dataset, 8 for CXR-8 and 5 for OpenI-5. We allocate 10% of the training data as validation set and conduct training over 5 epochs with a cosine decay learning rate schedule with linear warmup for 10% of the total training steps. Base learning rates are set to $2e^{-5}$ for the Pneumonia dataset and $1e^{-5}$ for both CXR-8 and OpenI-5. We employ binary cross entropy as the loss function and the model selection for testing is based on the epoch with the lowest validation loss.
0.A.4 Zero-shot Classification
Following the CLIP paper[36], we generate descriptive prompts for each label, mirroring the patterns found in the radiology reports of our pretraining data. For example, within the pneumonia detection task, a ‘normal’ X-ray is prompted as “Chest radiograph with normal findings, no signs of pneumonia”, while a prompt for pneumonia diagnosed X-rays would read “Radiograph of the chest displaying multifocal opacities, suggestive of viral pneumonia”. We apply the ensemble promoting technique, where we generate multiple variations of each label’s prompt to create a list of text embeddings for each label. The mean of these embeddings serves as the representation for the corresponding label. The specific prompts utilized for each label have been included in the Supplement section.
These prompts are converted into embeddings using the text encoder of the trained CLIP model, while the images are processed using the corresponding image encoder to produce image embeddings. Classification is then performed by selecting the label whose text emebdding is most similar to the image embedding, as determined by cosine similarity. We use the prompts used in [15], samples from which are shown below.
{mdframed}
[backgroundcolor=gray!10,linewidth=2pt] ZS Prompts
Atelectasis - mild subsegmental atelectasis
Cardiomegaly - cardiac silhouette size is mildly enlarged
Consolidation - increased reticular consolidation at the lower lung zone
Edema - mild pulmonary edema
Pleural Effusion - stable right bilateral pleural effusion
Pneumonia - Bronchopneumonia pattern suggestive of bacterial infection
Appendix 0.B Additional Results
In Table 6 we show the zero-shot classification accuracies for the CLIP, eCLIP and baseline models.
Table 6: Zero-shot classification performance on 4 X-ray datasets and model configurations, reported as accuracy from three independent random seeds. The highest score per dataset and model configuration is underlined. The overall best-performing model for each dataset is highlighted in bold.
Model Dataset
Chexpert 5x200 MIMIC 5x200 RSNA CXR 14x100
[…]
    ${}_{+\text{naive}}$ $0.409_{\pm.061}$ $0.369_{\pm.044}$ $0.669_{\pm.049}$ $0.145_{\pm.020}$
    ${}_{+\text{DACL}}$ $0.530_{\pm.016}$ $0.438_{\pm.007}$ $0.752_{\pm.007}$ $0.179_{\pm.008}$
[…]
    +expert (${}_{+\text{expert (ours)}}$ $0.518_{\pm.001}$ $0.436_{\pm.008}$ $0.730_{\pm.023}$ $0.179_{\pm.028}$
    ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ $0.520_{\pm.004}$ $0.478_{\pm.008}$ $0.753_{\pm.000}$ $0.168_{\pm.001}$
CLIPSwin ${}_{\text{Swin Tiny}}$ $0.529_{\pm.020}$ $0.454_{\pm.003}$ $0.799_{\pm.004}$ $0.188_{\pm.006}$
    ${}_{+\text{naive}}$ $0.536_{\pm.008}$ $0.459_{\pm.018}$ $0.793_{\pm.002}$ $0.188_{\pm.012}$
    ${}_{+\text{DACL}}$ $0.481_{\pm.005}$ $0.398_{\pm.020}$ $0.758_{\pm.005}$ $0.124_{\pm.019}$
[…]
    +expert (${}_{+\text{expert (ours)}}$ $0.549_{\pm.016}$ $0.443_{\pm.023}$ $0.799_{\pm.004}$ $0.195_{\pm.003}$
    ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ $0.565_{\pm.004}$ $0.465_{\pm.006}$ $0.810_{\pm.001}$ $0.210_{\pm.001}$
CLIPViT ${}_{\text{ViT Small}}$ $0.524_{\pm.023}$ $0.441_{\pm.007}$ $0.796_{\pm.000}$ $0.179_{\pm.007}$
    ${}_{+\text{naive}}$ $0.532_{\pm.017}$ $0.454_{\pm.028}$ $0.793_{\pm.003}$ $0.169_{\pm.022}$
    ${}_{+\text{DACL}}$ $0.485_{\pm.024}$ $0.402_{\pm 015}$ $0.753_{\pm.009}$ $0.154_{\pm.007}$
[…]
    +expert (${}_{+\text{expert (ours)}}$ $0.548_{\pm.017}$ $0.456_{\pm.010}$ $0.795_{\pm.005}$ $0.185_{\pm.015}$
    ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ $0.556_{\pm.001}$ $0.437_{\pm.004}$ $0.805_{\pm.001}$ $0.180_{\pm.003}$
CLIPViT ${}_{\text{ViT Base}}$ $0.540_{\pm.011}$ $0.470_{\pm.007}$ $0.788_{\pm.012}$ $0.200_{\pm.006}$
    ${}_{+\text{naive}}$ $0.503_{\pm.010}$ $0.434_{\pm.011}$ $0.787_{\pm.002}$ $0.176_{\pm.008}$
    ${}_{+\text{DACL}}$ $0.484_{\pm.006}$ $0.406_{\pm.002}$ $0.737_{\pm.001}$ $0.204_{\pm.098}$
[…]
    +expert (${}_{+\text{expert (ours)}}$ $0.564_{\pm.023}$ $0.476_{\pm.005}$ $0.799_{\pm.003}$ $0.204_{\pm.020}$
    ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ $0.557_{\pm.011}$ $0.464_{\pm.018}$ $0.782_{\pm.010}$ $0.197_{\pm.0.013}$
Figure 6 demonstrates the sample efficiency of different models. The top row shows the zero-shot performance on three multi-label classification test sets for DACL and eCLIP Swin Tiny models, trained with varying amounts of training batches. This highlights each model’s ability to generalize with limited data. The bottom row presents the linear probe scores for m3-mixup and eCLIP Swin Tiny models, evaluated with different amounts of training data. This illustrates how quickly each model learns and performs as more data is provided. These results underscore the importance of sample efficiency in model performance.
Figure 6: Sample Efficiency. (top row) Zero-shot performance on three multi-label classification test sets for DACL and eCLIP Swin Tiny models, trained with varying amounts of training batches. (bottom row) Linear probe scores with varying amounts of training data for $m^{3}$-mixup and eCLIP Swin Tiny models.
Appendix 0.C Visualize Embeddings with UMAP
Figure 7: 2D UMAP Projection of Embeddings Figure shows the UMAP projection of the Image, Text and heatmap processed Image embedding generated by eCLIP with Swin Tiny encoder. We use Open-I dataset for image and text and since expert annotation is unavailable for this dataset, we generate random uniform masks to simulate heatmaps.
We utilize the trained eCLIP model with the Swin Tiny encoder to examine UMAP projections of embeddings derived from the Open-I dataset. This dataset is categorized into subgroups based on the presence of specific abnormalities, as indicated in the ‘Problem’ column, which contains radiologist annotations for each image. Four primary abnormalities –‘normal,’ ‘cardiomegaly,’ ‘atelectasis,’ and ‘opacity’ – form the basis of our subgroup categorization. We ensure that samples within each subgroup are mutually exclusive, containing only one of these abnormalities.
To generate embeddings, we use the trained image and text encoders from our eCLIP model. Since the Open-I dataset lacks actual expert-annotated heatmaps, we simulate this condition by creating random heatmaps for each image. Thus we obtain the standard image embedding ($v_{i}$), text embedding ($t_{i}$) and expert image embedding ($v_{i}^{E})$ for each sample across the subgroups. These embeddings are projected into a 2D space using UMAP with cosine similarity as the metric. Subsequently, we visualize the 2D UMAP projections for each subgroup separately, facilitating a detailed inspection of the embedding distribution and the influence of expert annotations on the model’s representation space. This is shown in Fig. 7
Appendix 0.D Retrieval Augmented Generation of Radiologist Report
Figure 8: Retrieval Augmented Generation of Radiology Reports. Radiology reports from the training corpus are encoded using the CLIP/eCLIP’s text encoder to obtain text embeddings, which are then stored in a FAISS vector database. For a test image, the corresponding image embedding is obtained using the CLIP/eCLIP’s image encoder. This image embedding is queried against the FAISS database to find the nearest text embeddings, which are used as prompts for the Mistral 7B Large Language Model (LLM). The LLM then generates a test report based on these prompts.
We detail our approach to generate radiologist reports by augmenting a Large Language Model (LLM) with retrieved report snippets from the training corpus. This method is designed to tackle the challenge of generating medically relevant and coherent radiology report without direct image examination or explicit fine-tuning on medical datasets. The process involves the following key steps:
  1. 1.
    Text Embedding and Indexing: We utilize the Open-I dataset to create the source of the radiology reports. These reports are processed through the CLIP/eCLIP trained text encoder to produce embeddings of dimension 512. These embeddings are then normalized and indexed using FAISS [8] vector database, facilitating efficient retrieval based on similarity.
  2. 2.
    Text Retrieval and Clustering: For a given test X-ray image, we first compute its embedding using the CLIP/eCLIP trained image encoder and then query the FAISS index to retrieve the closest report snippets. The “closeness” is based on cosine similarity of the normalized embeddings, ensuring that the retrieved texts are semantically relevant to the image’s medical context. We use K-means to categorize the embeddings of these snippets into distinct groups. This clustering ensures the selection of representative sentences that encapsulate the primary observations within each group, thereby preserving the diversity of the retrieved reports. We require five closest snippets for prompting the LLM, so we retrieve four times this from the FAISS index for clustering.
  3. 3.
    Report Generation with LLM: Leveraging the retrieved snippets as context, we employ a frozen LLM, Mistral 7B[19], to generate a comprehensive radiology report. The aim is to produce reports that closely mimic those written by radiologists, based on the insights from the retrieved texts.
Fig. 8 shows the schematic of the steps involved in report generation using a frozen LLM.
0.D.0.1 LLM Prompts
We use the following system and user prompts to use the frozen Mistral 7B model to generate the radiology report. We provide two exemplars of retrieved reports and the corresponding generated report as samples for the In-Context Learning (ICL) as the user prompt for the LLM.
{mdframed}
[backgroundcolor=gray!10,linewidth=2pt] <SYSTEM PROMPT>
You are to act as a radiologist, trained to generate radiology reports. Your task is to synthesize the information from the closest report snippets provided below into a comprehensive and medically accurate radiologist report for each case. Craft a comprehensive response that is concise, succinct, and focuses on the key findings and potential diagnoses. Your report should maintain a professional tone, with clarity and precision in medical terminology, suitable for medical experts. Remember to be concise, succinct, and focus on the key findings and potential diagnoses, avoiding unnecessary elaboration.
{mdframed}
[backgroundcolor=gray!10,linewidth=2pt] USER
The following snippets are from reports closely related to the patient’s X-ray image.
< Retrieved Text >
Based on these, generate a radiologist report.
0.D.0.2 Evaluation Metrics
For evaluating the generated radiology reports, we employ metrics traditionally used in text generation and translation fields, namely BLEU-2 and BERT Score [64]. We also compute the cosine similarity between the embeddings of the generated reports and the ground truth, as derived from a reference model. This approach allows for a broader assessment of semantic congruence. Specifically, we utilize sentence transformer models [38] known for their effectiveness in sentence-level comparison tasks. We employ the ‘all-mpnet-base-v2’ [42] model for its general semantic understanding, and the ‘CheXBERT’ model [41], for its domain-specific performance in medical classification tasks. These models facilitate a more comprehensive and contextually relevant evaluation of the linguistic and clinical content of the generated reports.
0.D.0.3 Generated Radiology Report Samples
We provide more randomly sampled generated radiology reports in Tab. 7
Table 7: More Random samples of generated report. For each image in the Open-I dataset, the five closest text snippets based on embedding cosine similarity is used as prompts for Mistral 7B LLM. Utilizing in-context learning, we prompt the LLM with two such snippet-report pairs.
Problems Ground Truth Generated
normal The heart is normal in size. The mediastinum is unremarkable. The lungs are clear.. The lungs are clear with no evidence of consolidation, pleural effusion, or pneumothorax. Heart size is within normal limits. The visualized osseous structures are unremarkable.
normal XXXX XXXX and lateral chest examination was obtained. One AP view is expiratory and was repeated. The heart silhouette is normal in size and contour. Aortic XXXX appear unremarkable. Lungs demonstrate no focal infiltrates. There is no effusion or pneumothorax. The heart size is normal. The lungs are clear with no evidence of consolidation, pneumothorax, or pleural effusion. No acute abnormalities are noted in the visualized osseous structures. The previously seen right upper lobe mass lesion is not present on the current study.
Pulmonary Atelectasis; Pleural Effusion; Cardiomegaly; Pulmonary Congestion AP and lateral views were obtained. Bibasilar atelectasis and small left-sided pleural effusion. Stable cardiomegaly. No pneumothorax. Mild pulmonary vascular congestion. The lungs show bibasilar airspace opacities, possibly due to infection or atelectasis. Heart size is mildly enlarged. No pneumothorax or pleural effusion.
Lung; Density; Emphysema There is hyperexpansion. The heart size is normal. There is no pleural effusion or pneumothorax. Two circular densities overlying the right ribs which were not present in the XXXX CT. No focal infiltrates The lungs are hyperexpanded with clear fields and no acute consolidations or pleural effusions. Heart size and pulmonary vascularity are within normal limits. Degenerative changes are present in the thoracic spine. A calcified granuloma is noted in the left perihilar region. No pneumothorax or large pleural effusion is observed.
Lung; Spine Lungs are mildly hyperexpanded. The lungs are clear. There is no focal airspace consolidation. No pleural effusion or pneumothorax. Heart size and mediastinal contour are within normal limits. There are diffuse degenerative changes of the spine. The heart size and mediastinal contours are within normal limits. The lungs are clear, but with slightly low lung volumes and subtle left basal opacities. No pleural effusion, pneumothorax, or acute osseous abnormalities are noted. There is evidence of degenerative changes in the thoracic spine on the lateral view.
Calcified Granuloma; Opacity Heart size and pulmonary vascularity appear within normal limits. Calcified granuloma is present in the right base. No pneumothorax or pleural effusion is seen. In the lateral right base is identified an ill-defined somewhat oblong opacity. This was not present on the previous study. The remainder of the lungs appear clear. The lungs are clear with mild hyperexpansion, suggestive of underlying emphysema. Heart size is normal. No pleural effusions, pneumothoraces, or focal consolidations. The cardiomediastinal silhouette and osseous structures are unremarkable.